OpenStudy (anonymous):

Find the measure of an interior and an exterior angle of a regular 46-gon

OpenStudy (zale101):

Sum of interior angles = (n−2)∗180/n n is 46-gon You want to measure the 46-gon interior angles So = (46 - 2) * 180 / 46 (46 - 2) * 180 =7920 degrees Divide the interior measure 7920 by the n-gon which is 46 7920/46=172 ----------------------------- As for the exterior angles measure The sum of the exterior angles is 360/n And n=46 So 360/46=7.8 close to 8

OpenStudy (zale101):

http://www.mathwarehouse.com/geometry/polygon/ ~this site is helpful